OpenStudy (anonymous):

how to integrate 3y^(2)e^(xy) with respect to y?

OpenStudy (amistre64):

ideally, id say with integration by parts

OpenStudy (amistre64):

e^xy y^2 e^xy/x -2y e^xy/x^2 2 e^xy/x^3

OpenStudy (amistre64):

\[\frac{3}{x}e^{xy}(y^2-\frac{2}{x}y+\frac{2}{x^2})\]

OpenStudy (amistre64):

i spose that might be prettier if i factor out the unders all at once like the wolf would do

OpenStudy (amistre64):

\[\frac{3}{x^3}e^{xy}(x^2y^2-2xy+2)\]